Savitri has a sum of Rs x. She spent Rs 1000 on grocery, Rs 500 on clothes and Rs 400 on education, and received Rs 200 as a gift. How much money (in Rs) is left with her?
(A) x – 1700 (B) x – 1900 (C) x + 200 (D) x – 2100

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
Option (A) is correct here.
From the question it is given that,
Savitri has a sum of Rs x
She spent money on grocery =₹1000
She spent money on clothes=₹500
She spent money on education=₹400
She received gift =₹200

Total money spent by Savitri = 1000 + 500 + 400 =₹1900
Then,
Total money left with her after deducting =₹(x - 1900)
Therefore, money left with her after adding gift money = (x - 1900) + 200

= x–1700




Thanks